OpenStudy (babynini):

Let x =2sin(theta), -pi/2< theta

OpenStudy (babynini):

\[\frac{ x }{ \sqrt{4-x^2} }\]

OpenStudy (babynini):

\[\frac{ 2\sin(\theta) }{ \sqrt{4-(2\sin(\theta))^2} }\]

OpenStudy (babynini):

@UsukiDoll not sure where to go with this one! \[\frac{ 2\sin(\theta) }{ 2-2\sin(\theta) }\] ??

OpenStudy (usukidoll):

hmm try factoring the 2 out of the denominator

OpenStudy (babynini):

er..how do I do this xP

OpenStudy (babynini):

is it 2(1-sintheta) ?

OpenStudy (babynini):

in the denominator of course.

OpenStudy (usukidoll):

\[\frac{ 2\sin(\theta) }{ 2(1-\sin(\theta)) }\] then cancel the 2

OpenStudy (babynini):

and we're left with (sin(theta))/(1-sin(theta))

OpenStudy (usukidoll):

and then it turns into a mess... geez...

OpenStudy (babynini):

haha aii

OpenStudy (usukidoll):

I'm wondering if that's it... It's been a while. but I know that the domain is restricted counter clockwise 90 degrees to clockwise 90 degrees.

OpenStudy (babynini):

what if we ^2 the whole thing?

OpenStudy (usukidoll):

wait...

OpenStudy (babynini):

we can add 180 later.. right? to get it into the correct domain.

OpenStudy (usukidoll):

\[\frac{ 2\sin(\theta) }{ \sqrt{4-(2\sin(\theta))^2} }\] \[\frac{ 2\sin(\theta) }{ \sqrt{4-4\sin^2(\theta))} }\] \[\[\frac{ 2\sin(\theta) }{ \sqrt{4(1-\sin^2(\theta))} }\]\]

OpenStudy (usukidoll):

\[\frac{ 2\sin(\theta) }{ \sqrt{4(1-\sin^2(\theta))} }\]

OpenStudy (usukidoll):

\[\cos^2x+\sin^2x=1 \] \[\cos^2x = 1-\sin^2x\] \[\frac{ 2\sin(\theta) }{ \sqrt{4(\cos^2(\theta))} }\]

OpenStudy (usukidoll):

\[\frac{2\sin(\theta)}{2\cos(\theta)}\]

zepdrix (zepdrix):

\[\Large\rm \sqrt{4-4\sin^2x}\ne 2-2\sin x\]You silly billy Miriam -_-

OpenStudy (usukidoll):

I got tangent theta in return?!

OpenStudy (usukidoll):

I saw something weird when I saw that I mean can't we yank the 4 out and use a trig identity

OpenStudy (babynini):

Woah how'd you get tangent out of there? because the sin and cos are the same here?

OpenStudy (usukidoll):

\[\sqrt{4-4\sin^2x} \rightarrow \sqrt{4(1-\sin^2x)}\]

OpenStudy (usukidoll):

\[\sqrt{4\cos^2x} \rightarrow 2cosx \]

OpenStudy (usukidoll):

\[\frac{2sinx}{2cosx} \rightarrow \frac{sinx}{cosx} \rightarrow tanx\]

OpenStudy (babynini):

oou..I see!

OpenStudy (usukidoll):

you've skipped a step and that trickled down later on

OpenStudy (usukidoll):

\[(2\sin(\theta))^2 \rightarrow (2\sin(\theta))(2\sin(\theta))\]

OpenStudy (usukidoll):

\[4\sin^2(\theta)\]

OpenStudy (usukidoll):

you've neglected to take the 2sin(theta) to the second power... that's why everything became nonsense

OpenStudy (babynini):

Sorry, my internets really bad. yeah I was jumping around too much without checking o.0

OpenStudy (babynini):

Sowriee